If x^3 < 0, then sqrt (25x^2) +5x= ?
The answer is supposed to be 0, but I don't see how they got that.

Answer by Boreal(15235)

sqrt(25x^2)+5x=
x^3<0, so x must be less than 0.
sqrt(25x^2) is normally +/-5x. It can be +5x or-5x.
BUT we know x<0.
Therefore, the sqrt (25x^2)= MINUS 5x ONLY.
-5x+5x=0.